OpenStudy (anonymous):

Let f be a function defined for x>0 such that f'(x)=1/x for all x>o and f(1)=0 show that f(ab)=f(a)+f(b), f(a/b)=f(a)-f(b) and f(a^r)=rf(a) for all a,b>0, r (do not use f(x)=lnx)

OpenStudy (anonymous):

Could not integrate f'(x) then prove?

OpenStudy (anonymous):

could we use lnx+c

OpenStudy (anonymous):

I would use that too if it is possible, @anonymoustwo44

OpenStudy (anonymous):

if f'(x)=1/x then f(x) could only be \[\ln \left| x \right|\]

OpenStudy (anonymous):

f'(x) =1/x f(x) = ln| x |+C f(1)=0 ln1+C=0 C=0 f(x)=ln| x| f(ab) = ln|ab|=ln |a| +ln|b| = f(a)+f(b) f(a/b)=ln |a/b|=ln|a|-ln|b| =f(a)-f(b) f(a^r)=ln|a^r|=rln|a|=rf(a)

OpenStudy (anonymous):

Will you do it like this, @anonymoustwo44

OpenStudy (anonymous):

yes exactly :))